Please help me on this proportion.
A 60 meter steel pole is cut into two parts in the ratio of 11:4. How much longer is the longer part than the shorter?
Thank you

Answer by AnlytcPhil(1116)

Please help me on this proportion.
A 60 meter steel pole is cut into two parts in the ratio of 11:4. 
How much longer is the longer part than the shorter?

Let x meters = the length of the longer piece.

Then the length of the shorter piece is 
60 meters - x meters or 
60 - x meters

  longer piece     11
——————————————— = ————
 shorter piece      4

    x       11
———————— = ————
 60 - x      4

Cross multiply:

(LEFT NUMERATOR)×(RIGHT DENOMINATOR) =
              (RIGHT NUMERATOR)×(LEFT DENOMINATOR)

(x)(4) = (11)(60 - x)

    4x = 11(60 - x)

    4x = 660 - 11x

   15x = 660

     x = 660/15

     x = 44 = length of longer piece

length of shorter piece = 60 - x = 60 - 44 =
16 meters

To find how much longer 44 is than 16, we subtract
44 - 16 and get 28 meters.

Edwin
AnlytcPhil@aol.com